OpenStudy (anonymous):

Solve using the elimination method: 7x/2 + 9y/2=1 and x+7y/6=1

hero (hero):

Multiply The first equation by 2; Multiply the second equation by 6. 7x + 9y = 2 6x + 7y = 6 Solve both equations for y: y = (2-7x)/9 y = (6-6x)/7 Set y = y (by elimination method): (2-7x)/9 = (6-6x)/7 Cross Multiply: 9(6-6x) = 7(2-7x) Distribute: 54 - 54x = 14 - 49x Continue Solving for x: 6x = 40 x = 40/6 or 20/3 Substitute x back in to the original equation to find y